OpenStudy (anonymous):

Solve this system algebraically. 9x + 2y = 5 y - 2x + 3 = 0 {(11/13, -17/13)} {(-11/13, -17/13)} {(-11/13, 17/13)} {(11/13, 17/13)}

OpenStudy (ttp):

(x=11/13, y=-17/13) It's A. Medal?